4- One number is four more than a second number. Two times the first number is 8 more than four times the second number. Find the numbers?

Answer by TimothyLamb(4379)

x = y + 4
2x = 4y + 8
---
put the system of linear equations into standard form
---
x - y = 4
2x - 4y = 8
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 4
y = 0
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php